### v2

@@ 4 Structural Interpretation of Retrospective Analysis
We begin with a demand function $Q=D(P,Z;\boldsymbol{\alpha})+\varphi$, where $Q$ is the quantity demanded of the homogenous good, $P$ is its price, $Z$ is a random variable that captures stochastic factor that shifts demand and is not affected by $P$ or $\varphi$, and $\boldsymbol{\alpha}$ is the vector of demand parameter(s). To be consistent with the reduced form (1), we maintain the assumption of homogeneous demand. The corresponding marginal revenue, augmented to allow for the conduct, is given by $MR=P+\lambda\times h(\cdot)$, where $h(Q)=P^{\prime}(Q)\times Q\leq 0$ and $\lambda\in[0,1]$ is the strategic conduct parameter. The conduct parameter $\lambda$ measures the degree of competition in the market, where $\lambda=0$ represents perfect competition, and $\lambda=1$ represents monopolistic behavior.
Let $c(Q,W,I;\boldsymbol{\theta})$ denote the marginal cost of production, where $Q$ is the quantity supplied of the homogenous good, $W$ is a random variable that captures stochastic factor that shifts cost and is not affected by $P$ or $\varphi$, $I\in\{0,1\}$ is an indicator variable that is equal to one post-merger and zero otherwise, and $\boldsymbol{\theta}$ is the vector of cost parameter(s).
@@
A key feature of our specification is that the merger affects only the supply relationship—specifically, the cost and conduct parameters—rather than the demand function itself. However, this feature does not mean the merger will not affect demand-side outcomes such as prices. Rather, it assumes that preference parameters (like consumer price sensitivity) remain unchanged by the merger. There is no reason why consumers’ preferences should change just because fewer firms now offer a homogeneous product; consumers only respond to the resulting price changes.
Furthermore, our specification also allows for various post-merger scenarios, from maintained competition, i.e., when $\lambda(1)\approx\lambda(0)$, to enhanced market power, i.e., $\lambda(1)>\lambda(0)$, and merger induced efficiency gain, i.e., $c(\cdot,\cdot,1;\cdot)\leq c(\cdot,\cdot,0;\cdot)$.
@@
where the variable $I_{mt}$ equals 0 for control markets across all periods $t$ and equals 1 for the treated markets in periods after the merger. Therefore, we can interpret $I_{mt}=\text{Treatment}_{mt}\times\text{Post-Merger}_{mt}$ as the main variable of interest and its coefficient $\beta^{\dagger}:=\frac{\theta_{3}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}$ as the effect of merger on prices.
The intercept in (5) corresponds to the combined effect of the constant term plus market and time-fixed effects in Equation (1). Additionally, the exogenous variable $X_{mt}$ in (1) represents the vector of demand and cost shifters $(Z_{mt},W_{mt})$. Now, we can compare (5) with the DiD specification (1) to understand the structural interpretation of the reduced form regressions. [^11] The dependent variables in Equation (1) are log prices, while they are in levels in Equation (5). We refer the reader to Appendix Table B.1 Panel B for proper comparison. The marginal effects are analogous whether we use levels or logs. For a streamlined presentation, we present the log specification in the main text and the level specification in the appendix.
@@
Second, the composite error ($\varkappa_{mt}$) in Equation (1) contains both demand ($\varphi$) and supply ($\vartheta$) shocks. So, even if merger decisions are independent of market conditions and firms conduct does not change (i.e., $\lambda(1)=\lambda(0)$), consistent estimation requires that mergers neither alter firm conduct nor experience differential evolution of unobservables across treatment and control markets, i.e., $\mathbb{E}(I_{mt}\mid\varkappa_{mt})=0$.
Third, while cost shocks ($\vartheta_{mt}$) may plausibly be uncorrelated with merger decisions, demand shocks ($\varphi_{mt}$) could systematically differ between merging and non-merging markets —particularly if firms with declining demand are more likely to merge. [^12] Interestingly, both the AA-TWA merger involved TWA’s hub in St. Louis, a city that had been declining in economic importance, and the UA-CO merger involved Cleveland, another city facing economic challenges. While we do not establish a causal relationship, these examples suggest the possibility of such patterns on the demand side.
If mergers affect firms’ conduct, the DiD estimates may be biased because the DiD specification fails to account for an indirect effect. This effects operates through the composite error $\varkappa_{mt}$ via the relationship $\tilde{\gamma}_{mt}=\theta_{1}-\lambda(I_{mt})/\alpha_{1}$. Specifically, if a merger changes firms’ conduct, it alters $\varkappa_{mt}$, creating a correlation between the merger indicator $I_{mt}$ and the error term $\varepsilon_{mt}$. This correlation, i.e., $\mathbb{E}(I_{mt}\mid\varkappa_{mt})\neq 0$, violates a key assumption of DiD estimation, resulting in biased estimates.
Crucially, even when these identification assumptions fail, and the reduced-form DiD estimates become biased, our structural framework maintains its identification. This key advantage underscores the importance of complementing DiD analysis with structural approaches when evaluating mergers.
@@
Building on the discussion of conduct parameter identification, we can draw a meaningful parallel to Miller and Weinberg (2017) who consider differentiated products. Despite differences in our models, they also estimate mergers’ effect on conduct using pre- and post-merger data. Such before-and-after comparison is essential for identifying the coordinated effect (Berry and Haile, 2014), or we can use simulation like in Igami and Sugaya (2022).
For estimation, we use several exogenous variables, not just one. On the demand side, we use market size (defined as the geometric mean of endpoint populations) and demographic characteristics at both origin and destination cities: per capita income, net migration, population size, and births and deaths. On the supply side, we use the nonstop flight ratio and year-quarter dummies to control for seasonality. These variables are listed with their summary statistics Appendix Table A.1131313We use the generalized method of moments estimator (Hansen, 1982). For inference, we estimate the Eicker-White-Huber heteroskedasticity-consistent standard errors (Wooldridge, 2010).
Table 5 presents estimates from our structural model across three specifications and the implied elasticities and reduced-form merger price effects. Panel A reports key structural parameters: the price coefficient ($\alpha_{1}$) capturing consumer price sensitivity, the merger’s effect on costs ($\theta_{3}$), and changes in the conduct parameter ($\Delta\lambda$). For market structure-specific
Table 5: Conduct Parameter Estimates and Implied Elasticities
@@
Price Coefficient ($\alpha_{1}$) -0.188∗∗∗ -0.027∗∗∗ -0.092 -0.188∗∗∗ -0.03∗∗∗ -0.092∗∗∗ -0.188∗∗∗ -0.03∗∗∗ -0.085∗∗∗
(1.66e-05) (8.59e-06) (0.1366) (1.05e-05) (9.2e-11) (1.2e-08) (1.5e-09) (1.03e-10) (1.8e-08)
Merger’s effects on Cost ($\theta_{3}$) 0.362∗∗∗ -0.667∗∗∗ 0.251∗∗∗ -0.134∗∗∗ 0.069∗∗∗ -0.059∗∗∗ -0.134∗∗∗ 0.069∗∗∗ 0.059∗∗∗
(0.0001) (0.0001) (0.0075) (9.878e-05) (3.2e-13) (2.7e-08) (3.4e-08) (7.1e-14) (4.05e-08)
@@
(0.0001) (2.03e-10) (2.3e-08)
“2 to 1” markets ($\Delta\lambda_{2}$) 0.161∗∗∗ 0.023∗∗∗ 0.076∗∗∗
(7.2e-09) (6.3e-11) (2.5e-08)
“3 to 2” markets ($\Delta\lambda_{3}$) 0.187∗∗∗ 0.027∗∗∗ 0.084∗∗∗
(1.07e-08) (1.3e-10) (3.6e-09)
@@
Mean Elasticity -0.671 -0.095 -0.253 -0.670 -.0103 -0.252 -0.670 -0.103 -0.235
Median Elasticity -0.128 -0.020 -0.075 -0.128 -0.021 -0.074 -0.128 -0.021 -0.069
@@
“2 to 1” markets -0.125∗∗∗ 0.068∗∗∗ 0.057∗∗∗
(3.1e-08) (3.9e-12) (3.9e-08)
“3 to 2” markets -0.122∗∗∗ 0.068∗∗∗ 0.057∗∗∗
(3.1e-08) (3.5e-12) (3.8e-08)
  • Notes: Panel A presents GMM estimates of price coefficient ($\alpha_{1}$) and merger coefficient in costs ($\theta_{3}$) in the structural parameters from Equation (4). Column (1) imposes competitive conduct ($\lambda=0$), Column (2) allows for a homogeneous conduct parameter that differs only between pre- and post-merger periods, and Column (3) allows conduct to vary by market structure. The changes in conduct are the estimates of the parameter defined in Equation (6). Panel B presents the implied mean and median elasticities, and Panel C presents the implied effect of the merger on prices in Equation (5). Eicker-White-Huber heteroskedasticity-consistent standard errors are reported in parentheses. ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$.
conduct in column (3), we distinguish between the 3 to 2 markets ($\Delta\lambda_{3}$) and 2 to 1 markets ($\Delta\lambda_{2}$), measuring changes in competitive behavior following the mergers.
@@
Our preferred specification in Column (3) provides a more nuanced analysis by allowing differential conduct changes in “2 to 1” markets ($\Delta\lambda_{2}$) versus “3 to 2” markets ($\Delta\lambda_{3}$). This market structure-specific approach reveals that both types of markets experienced statistically significant increases in market power post-merger. Interestingly, we find that “3 to 2” markets generally saw larger increases in the conduct parameter than “2 to 1” markets across all three mergers. For instance, in the DL-NW merger, $\Delta\lambda_{3}$ (0.187) exceeds $\Delta\lambda_{2}$ (0.161), suggesting that markets with remaining competition still experienced substantial competitive harm.
Statistical tests strongly reject the null hypothesis of no change in conduct for all three mergers, confirming significant increases in market power. The magnitudes of the conduct parameter changes (ranging from 0.023 to 0.188) are smaller than the 0.37 estimate reported by Miller and Weinberg (2017) for the beer industry after a joint venture. This difference likely reflects the airline industry’s network structure, where the presence of connecting routes and alternative carriers maintains some competitive pressure even after consolidation. However, mergers need not always increase coordinated effects. Igami and Sugaya (2022) demonstrate that mergers can simultaneously reduce competitiveness (unilateral effects) while decreasing collusion (coordinated effects). These findings suggest that coordinated effects depend on specific industry characteristics, firm asymmetries, and mergers.
@@
Focusing on our preferred specification in Column (3), the implied reduced-form estimates reveal heterogeneous price effects across the three mergers. The DL-NW merger led to price reductions ($\beta^{\dagger}\approx-0.12$), suggesting that efficiency gains outweighed increased market power. In contrast, the UA-CO and AA-US mergers led to a modest price increase ($\beta^{\dagger}\approx 0.068$ and $0.057$, respectively).
While these results provide valuable insights into the relationship between market structure and firm conduct, they rest on assumptions of homogeneous products, linear demand, and linear costs. To assess the robustness of our findings, Section 6 extends our analysis using discrete choice models that relax both the homogeneity and linear demand assumptions, allowing for a more flexible characterization of consumer preferences and market competition. Before that, we show how to use only pre-merger data to run reduced form analysis.
@@ 5 A Reduced-Form Approach to Merger Prediction
We present a reduced-form approach that we can use to understand merger effects when only the pre-merger data are available. Specifically, we propose a method that fully leverages the assumption that market structures are exogenous. If market structure changes are exogenous, pre-merger transitions—from duopoly to monopoly or triopoly to duopoly—should inform us about post-merger outcomes in 2 to 1 and 3 to 2 markets. This methodology builds on Baker (1999)’s analysis of the Staples-Office Depot merger, later expanded by Davis (2005) to study movie theaters, and recently employed by Bruegge et al. (2024) using entry events to simulate the impact of the merger on prices.
@@
  • Notes: This table presents the coefficients from Equations (7) and (8), for each of the three mergers, estimated using first difference method. Panel A reports coefficients on simulated changes in HHI in Equation (7). Panel B shows coefficients for market structure transitions, where $2\rightarrow 1$ indicates transition from duopoly to monopoly, $1\rightarrow 2$ from monopoly to duopoly, and anologously for other variables as defined in Equation (8). All dependent variables (price, passengers, seats) are in logarithms using the Top 50 markets by population. Standard errors in parentheses. * p$<$0.10, ** p$<$0.05, *** p$<$0.01.
@@
where the indicator variables capture market structure transitions. For example, $\mathbbm{1}\{\text{Duopoly}\rightarrow\text{Monopoly}\}_{mt}$ equals 1 if market $m$ transitioned from duopoly to monopoly at time $t$.
@@
Our analysis of exogenous market structure transitions reveals consistent patterns across all three airline mergers. Consolidation leads to moderate price increases but substantial capacity reductions, while market entry produces symmetric effects in the opposite direction. Bruegge et al. (2024), in their investigation of the Jet Blue and Spirit Air merger, where they observed that “entries with larger capacity tend to have greater fare effects, and that the difference in fare effects between carriers varies with the capacity level.” Our analysis shows significant capacity changes but relatively modest price effects across the mergers we studied.

### v3

@@ 4 Structural Interpretation of Retrospective Analysis
We begin with a demand function $Q=D(P,Z;\boldsymbol{\alpha})+\varphi$, where $Q$ is the quantity demanded of the homogenous good, $P$ is its price, $Z$ is a random variable that captures the stochastic factor that shifts demand and is not affected by $P$ or $\varphi$, and $\boldsymbol{\alpha}$ is the vector of demand parameter(s). To be consistent with the reduced form (1), we maintain the assumption of homogeneous demand. The corresponding marginal revenue, augmented to allow for the conduct, is given by $MR=P+\lambda\times h(\cdot)$, where $h(Q)=P^{\prime}(Q)\times Q\leq 0$ and $\lambda\in[0,1]$ is the strategic conduct parameter. The conduct parameter $\lambda$ measures the degree of competition in the market, where $\lambda=0$ represents perfect competition, and $\lambda=1$ represents monopolistic behavior.
Let $c(Q,W,I;\boldsymbol{\theta})$ denote the marginal cost of production, where $Q$ is the quantity supplied of the homogenous good, $W$ is a random variable that captures the stochastic factor that shifts cost and is not affected by $P$ or $\varphi$, $I\in\{0,1\}$ is an indicator variable equal to one post-merger and zero otherwise, and $\boldsymbol{\theta}$ is the vector of cost parameter(s).
@@
A key feature of our specification is that the merger affects the supply relationship, specifically, the cost and conduct parameters, but not the demand function. However, this feature does not mean the merger will not affect demand-side outcomes such as prices and product characteristics. Rather, it assumes that preference parameters (like consumer price sensitivity) remain unaffected by the merger. There is no reason why consumers’ preferences should change just because fewer firms now offer a homogeneous product; consumers respond only to the resulting price changes and any changes in product characteristics, such as airlines’ networks. However, we continue to assume that, unlike the changes in prices, any change in product characteristics after a merger is exogenous. [^11] This is a widely used assumption in the literature, see, e.g., Ciliberto et al. (2021). More recently, Fan and Yang (Forthcoming) relaxed this assumption and allowed product characteristics to change following mergers.
Furthermore, our specification also allows for various post-merger scenarios, from maintained competition, i.e., when $\lambda(1)\approx\lambda(0)$, to enhanced market power, i.e., $\lambda(1)>\lambda(0)$, and merger-induced efficiency gain, i.e., $c(\cdot,\cdot,1;\cdot)\leq c(\cdot,\cdot,0;\cdot)$.
@@
where the variable $I_{mt}$ equals 0 for control markets across all periods $t$ and equals 1 for the treated markets in periods after the merger. So, $I_{mt}=\text{Treatment}_{mt}\times\text{Post-Merger}_{mt}$ is the main variable of interest and its coefficient $\beta^{\dagger}:=\frac{\theta_{3}}{(1-\tilde{\gamma}_{mt}\alpha_{1})}$ is the price-effect of merger.
The intercept in (5) corresponds to the combined effect of the constant term plus market and time-fixed-effects in Equation (1). Also, the exogenous variable $X_{mt}$ in (1) represents the vector of demand and cost shifters $(Z_{mt},W_{mt})$. Now, we can compare (5) with the DiD specification (1) to understand the structural interpretation of the reduced form regressions. [^12] The dependent variables in Equation (1) are log prices, while they are in levels in Equation (5). We refer the reader to Appendix Table B.1 Panel B for proper comparison. The marginal effects are analogous whether we use levels or logs. For a streamlined presentation, we present the log specification in the main text and the level specification in the appendix.
@@
Second, the composite error ($\varkappa_{mt}$) in Equation (1) contains both demand ($\varphi$) and supply ($\vartheta$) shocks. So, even if merger decisions are independent of market conditions and firms’ conduct does not change (i.e., $\lambda(1)=\lambda(0)$), consistent estimation requires that mergers neither alter firm conduct nor experience differential evolution of unobservables across treatment and control markets, i.e., $\mathbb{E}(I_{mt}\mid\varkappa_{mt})=0$.
Third, while cost shocks ($\vartheta_{mt}$) may plausibly be uncorrelated with merger decisions, demand shocks ($\varphi_{mt}$) could systematically differ between merging and non-merging markets —particularly if firms with declining demand are more likely to merge. [^13] Interestingly, the AA-TWA merger involved TWA’s hub in St. Louis, a city declining in economic importance, and the UA-CO merger involved Cleveland, another city facing economic challenges. While we do not establish a causal relationship, these examples suggest the possibility of such patterns in the demand.
If mergers affect firms’ conduct, the DiD estimates may be biased because the DiD specification fails to account for an indirect effect. Thi effects operates through the composite error $\varkappa_{mt}$ via the relationship $\tilde{\gamma}_{mt}=\theta_{1}-\lambda(I_{mt})/\alpha_{1}$. Specifically, if a merger changes firms’ conduct, it alters $\varkappa_{mt}$, creating a correlation between the merger indicator $I_{mt}$ and the error term $\varepsilon_{mt}$. This correlation, i.e., $\mathbb{E}(I_{mt}\mid\varkappa_{mt})\neq 0$, violates a key assumption of DiD estimation, resulting in biased estimates.
Crucially, even when these identification assumptions fail, and the reduced-form DiD estimates become biased, our structural framework maintains its identification. This key advantage underscores the importance of complementing DiD analysis with structural approaches when one is evaluating mergers.
@@
Building on the discussion of conduct parameter identification, we can draw a meaningful parallel to Miller and Weinberg (2017), who consider differentiated products. Despite differences in our models, they also estimate mergers’ effect on conduct using pre- and post-merger data. Such before-and-after comparison is essential for identifying the coordinated effect (Berry and Haile, 2014), or we can use simulation like in Igami and Sugaya (2022).
For estimation, we use several exogenous variables. On the demand side, we use market size (defined as the geometric mean of endpoint populations) and demographic characteristics at both origin and destination cities: per capita income, net migration, population size, and births and deaths. On the supply side, we use the nonstop flight ratio and year-quarter dummies to control for seasonality. These variables are listed in Appendix Table A.1.141414We use the generalized method of moments estimator (Hansen, 1982). For inference, we estimate the Eicker-White-Huber heteroskedasticity-consistent standard errors (Wooldridge, 2010).
Table 5 presents estimates from our structural model across three specifications and the implied elasticities and reduced-form merger price effects. Panel A reports the price coefficient ($\alpha_{1}$), the merger’s effect on costs ($\theta_{3}$), and changes in the conduct parameter ($\Delta\lambda$). For market structure-specific conduct (column (3)), we distinguish between the 3 to 2 ($\Delta\lambda_{3}$) and 2 to 1 market ($\Delta\lambda_{2}$), measuring changes in competitive behavior after the merger.
Table 5: Conduct Parameter Estimates and Implied Elasticities and Price Effects of Mergers
@@
Price Coefficient ($\alpha_{1}$) -0.188∗∗∗ -0.027∗∗∗ -0.092 -0.188∗∗∗ -0.03∗∗∗ -0.092∗∗∗ -0.188∗∗∗ -0.03∗∗∗ -0.090∗∗∗
(1.66e-05) (8.59e-06) (0.1366) (1.05e-05) (9.2e-11) (1.2e-08) (9.0e-12) (9.2e-11) (7.0e-08)
Merger’s Effects on Cost ($\theta_{3}$)
Uniform 0.362∗∗∗ -0.667∗∗∗ 0.251∗∗∗ -0.134∗∗∗ 0.069∗∗∗ -0.059∗∗∗
(0.0001) (0.0001) (0.0075) (9.878e-05) (3.2e-13) (2.7e-08)
‘2 to 1” markets -0.131∗∗∗ -0.008∗∗∗
(3.4e-08) (1.4e-07)
‘3 to 2” markets -0.132∗∗∗ 0.069∗∗∗ -0.024∗∗∗
(1.9e-11) (3.2e-13) (6.4e-08)
@@
(0.0001) (1.7e-15) (2.3e-08)
‘2 to 1” markets ($\Delta\lambda_{2}$) 0.164∗∗∗ 0.078∗∗∗
(1.6e-11) (6.7e-08)
‘3 to 2” markets ($\Delta\lambda_{3}$) 0.188∗∗∗ 0.03∗∗∗ 0.081∗∗∗
(5.3e-12) (1.7e-15) (6.5e-08)
@@
Mean Elasticity -0.671 -0.095 -0.253 -0.670 -0.103 -0.252 -0.670 -0.103 -0.231
Median Elasticity -0.128 -0.020 -0.075 -0.128 -0.021 -0.074 -0.128 -0.021 -0.067
@@
‘2 to 1” markets -0.122∗∗∗ -0.008∗∗∗
(1.8e-11) (1.3e-07)
‘3 to 2” markets -0.120∗∗∗ 0.068∗∗∗ -0.023∗∗∗
(1.8e-11) (6.5e-12) (1.3e-07)
  • Notes: Panel A presents GMM estimates of price (100s) coefficient ($\alpha_{1}$) and merger coefficient in costs ($\theta_{3}$) in the structural parameters (Equation (4), with prices in 100s and quantities in 10,000s). Column (1) imposes competitive conduct ($\lambda=0$), Column (2) allows for a homogeneous conduct parameter that changes pre- and post-merger (Equation (6)), and Column (3) allows conduct to vary by market structure. Panel B presents the implied mean and median elasticities, and Panel C presents the implied effect of the merger on prices (Equation (5)). The standard errors are reported in parentheses. ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$.
implied elasticities and reduced-form merger price effects. Panel A reports key structural parameters: the price coefficient ($\alpha_{1}$) capturing consumer price sensitivity, the merger’s effect on costs ($\theta_{3}$), and changes in the conduct parameter ($\Delta\lambda$). For market structure-specific conduct in column (3), we distinguish between the 3 to 2 markets ($\Delta\lambda_{3}$) and 2 to 1 markets ($\Delta\lambda_{2}$), measuring changes in competitive behavior following the mergers.
@@
Our preferred specification in Column (3) provides a more nuanced analysis by allowing differential cost and conduct changes in “2 to 1” markets ($\Delta\lambda_{2}$) and “3 to 2” markets ($\Delta\lambda_{3}$), for DL-NW and AA-US mergers and only “3 to 2” markets ($\Delta\lambda_{3}$) for UA-CO mergers because it has only “3 to 2” markets. This market structure-specific approach reveals that both types of markets experienced statistically significant increases in market power post-merger. Interestingly, we find that “3 to 2” markets generally saw larger increases in the conduct parameter than “2 to 1” markets across all three mergers. For instance, in the DL-NW merger, $\Delta\lambda_{3}$ (0.188) exceeds $\Delta\lambda_{2}$ (0.164), suggesting that markets with remaining competition still experienced substantial competitive harm.
Statistical tests strongly reject the null hypothesis of no change in conduct for all three mergers, confirming significant increases in market power. The magnitudes of the conduct parameter changes (ranging from 0.081 to 0.188) are smaller than the 0.37 estimate reported by Miller and Weinberg (2017) for the beer industry after a joint venture. This difference likely reflects the airline industry’s network structure, where the presence of connecting routes and alternative carriers maintains some competitive pressure even after consolidation. However, mergers need not always increase coordinated effects. Igami and Sugaya (2022) demonstrate that mergers can simultaneously reduce competitiveness (unilateral effects) while decreasing collusion (coordinated effects). These findings suggest that coordinated effects depend on specific industry characteristics, firm asymmetries, and mergers.
@@
Focusing on our preferred specification in Column (3), the implied reduced-form estimates reveal heterogeneous price effects across the three mergers. The DL-NW and AA-US mergers led to price reductions, $\beta^{\dagger}$ between $-0.120$ and $-0.122$, and $\beta^{\dagger}$ between $-0.008$ and $-0.023$, respectively, suggesting that efficiency gains outweighed increased market power. In contrast, the UA-CO merger led to a modest price increase ($\beta^{\dagger}\approx 0.068$).
While these results provide valuable insights into the relationship between market structure and firm conduct, they rest on assumptions of homogeneous products, linear demand, and linear costs. To assess the robustness of our findings, Section 6 extends our analysis using discrete choice models that relax both the homogeneity and linear demand assumptions, allowing for a more flexible characterization of consumer preferences and market competition. Before that, we show how to use only pre-merger data in a reduced-form approach to predict after-merger prices.
@@ 5 A Reduced-Form Approach to Merger Prediction
We present a reduced-form approach to determine merger effects when only the pre-merger data are available. Specifically, we propose a method that fully leverages the assumption that market structures are exogenous. If market structure changes are exogenous, pre-merger transitions—from duopoly to monopoly or triopoly to duopoly—should inform us about post-merger outcomes in 2 to 1 and 3 to 2 markets. This methodology builds on Baker (1999)’s analysis of the Staples-Office Depot merger, later expanded by Davis (2005) to study movie theaters, and recently employed by Bruegge et al. (2024) using entry events to simulate the impact of the merger on prices.
@@
  • Notes: This table presents the coefficients from Equations (7) and (8), for each of the three mergers, estimated using the first-difference method. Panel A reports coefficients on simulated changes in HHI in Equation (7). Panel B shows coefficients for market structure transitions, where $2\rightarrow 1$ indicates the transition from duopoly to monopoly, $1\rightarrow 2$ from monopoly to duopoly, and anologously for other variables as defined in Equation (8). All dependent variables (price, passengers, seats) are in logarithms using the Top 50 markets by population. Standard errors in parentheses. * p$<$0.10, ** p$<$0.05, *** p$<$0.01.
@@
where the indicator variables capture the change in market structures. For example, the variable $\mathbbm{1}\{\text{Duopoly}\rightarrow\text{Monopoly}\}_{mt}$ equals 1 if market $m$ transitioned from duopoly to monopoly at time $t$.
@@
Our analysis of exogenous market structure transitions reveals consistent patterns across all three airline mergers. Consolidation leads to moderate price increases but substantial capacity reductions, while market entry produces symmetric effects in the opposite direction. Bruegge et al. (2024), in their investigation of the Jet Blue and Spirit Air merger, where they observed that “entries with larger capacity tend to have greater fare effects, and that the difference in fare effects between carriers varies with the capacity level.” Our analysis shows significant capacity changes but relatively modest price effects across the mergers.